If a bank pays 15% compounded semi-annually, how much should be deposited now to have $1800 in 5 years from now?
Amount the needs to be deposited now = $

Found 3 solutions by mananth, math_tutor2020, ikleyn:
Answer by mananth(16946)

the formula for compound interest:
A = P(1 + r/n)^(nt)

Where,
P = principal amount
r = annual interest rate (15%)
n = number of times the interest is compounded per year (2)
t = time period (5 years)
A = amount at the end of the time period ($1800)
Substituting the values in the formula, we get:
$1800 = P(1 + 0.15/2)^(2*5)
$1800 = P(1 + 0.075)^10
$1800 = P(1.075)^10
P = $975.62 (rounded to two decimal places)

Answer by math_tutor2020(3817)


Answer: $873.35

Work Shown:
A = P*(1+r/n)^(n*t)
1800 = P*(1+0.15/2)^(2*5)
1800 = P*(1.075)^(10)
1800 = P*2.061032
P = 1800/2.061032
P = 873.348885
P = 873.35

Check:
A = P*(1+r/n)^(n*t)
A = 873.35*(1+0.15/2)^(2*5)
A = 1800.001915
A = 1800.00

Answer by ikleyn(52781)

.

To future generations of students who may read this post:


        Calculations by @mananth are  WRONG.

        Ignore his post,  for your safety.